Water enters the heater with 25 degree Celsius and leaves at 78 degree Celsius. The temperature change is Answer degree Fahrenheit.?


1
Expert's answer
2020-10-02T14:50:34-0400

Water enters at "T_1=25^oC,T_2=78^oC" is the temperature of leaving the water

Change in temperature= "\Delta T=T_2-T_1=78-25=53 ^oC"

So, we know the relation between temperature in degree Celsius and temperature in degree Fahrenheit


"T(^oF)= \frac{9}{5}(T^oC)+32"


"T(^oF)= \frac{9}{5}(53)+32"


"T(^oF)=127.4 ^oF"